OpenStudy (anonymous):

find f(x) domain \[f(x)=\log _{x-3}(x ^{2}-1)/(x ^{2}-4)\]

OpenStudy (amistre64):

a few issues that might influence domain are an x under the fraction the x in the base notation, dont think we can have log_0 and the x on the log argument itself; nothing 0 and under is good

OpenStudy (anonymous):

you mean \[x >3 \]

OpenStudy (anonymous):

?

OpenStudy (amistre64):

well, lets start with the log argument part: x^2 - 1 <= 0 is bad x^2 <= 1 x <= -1 , or x >=1 would be offlimits

OpenStudy (amistre64):

so we are defined by just that to be between -1 and 1

OpenStudy (amistre64):

nothing in the interval makes the bottom look bad so lets try to see if theres anything in there that would make the base part go bad

OpenStudy (amistre64):

there might be some issue in a -base for SOME numbers

OpenStudy (amistre64):

for instance: log-2(8) = y 8 = (-2)^y is not defined

OpenStudy (anonymous):

the domain is this \[(3,4)\cup(4,+\infty)\]

OpenStudy (anonymous):

amir.sat is right

OpenStudy (amistre64):

i messed up on my x^2 - 1 dint i :)

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

amir you defeated the 100level guy

OpenStudy (amistre64):

ive been defeated plenty of times ;)

OpenStudy (amistre64):

wolfram would have told me if I was wrong in the end im sure of it lol

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

but i've done that with my self if you have a lot of time i'll show you

OpenStudy (anonymous):

amir don't you want to give a question MedaL

OpenStudy (anonymous):

oww yeah

OpenStudy (anonymous):

\[{((-3 + x) (-1 + x^2))/(-4 + x^2)} \]